OpenStudy (anonymous):

FAN AND MEDAL A box has 1 red marble, 1 blue marble, and 4 green marbles of the same shape and size. Dan draws a marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing two green marbles in a row? Explain your answer.

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Queli

OpenStudy (anonymous):

@OM14forever

OpenStudy (anonymous):

Wouldn't it be 4/6 * 4/6=16/36=4/9

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

I think so. @LynFran Is that right?

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

50%

OpenStudy (mathmate):

@OM14forever Agree! @Ninjaslice01_GT

OpenStudy (anonymous):

show ur work